Problem 2-17

Two types of cars (Deluxe and Limited) were produced by a car manufacturer last year. Quantities sold, price per unit, and labor hours follow.

QUANTITY $/UNIT   
  Deluxe car 2,850 units sold $ 8,700 /car
  Limited car 6,350 units sold $ 10,200 /car
  Labor, Deluxe 21,750 hours $ 11 /hour
  Labor, Limited 29,555 hours $ 14 /hour

What is the labor productivity for each car? Provide two sets of figures: units per labor hour, and dollar of output per dollar of labor expense. (Round your answers to 2 decimal places.)

Labor Productivity

   units/hour     dollars
  Deluxe car          
  Limited car          
Answer #1

For Deluxe car:

Labor productivity (units/ hour) = 2850/21750 = .13 units per labor hour

Labor productivity (Dollars) = 2850*8700/(21750*11) = $103.64 output / dollar of labor expenses

For Limited car:

Labor productivity (units/ hour) = 6350/29555 = .21 units per labor hour

Labor productivity (Dollars) = 6350*10200/(29555*14) = $ 156.54 output / dollar of labor expenses

Now, following matrix is prepared on the basis of above calculation.


Labor Productivity

   units/hour

    dollars

  Deluxe car

.13    

$103.64   

  Limited car

.21    

$156.54